OpenStudy (anonymous):

can someone please help me prove the identity: cos^4 x- sin^4 x =cos^2 x- sin^2 x

OpenStudy (campbell_st):

well factor the left side \[(\cos^2(x) - \sin^2(x))(\cos^2(x) + \sin^2(x)) = (\cos^2(x) - \sin^2(x)) \] and you know \[\cos^2(x) + \sin^2(x) = 1\] so you should be able to prove it from here

OpenStudy (anonymous):

now i'm embarrassed, i totally forgot about factoring, thank you